=== 2019/7/24(Wed.) ===

==== Schedule ====
 * ~7/31: Trial for DRMI with polarization.
 * 7/6~: Trial for mass lock

 * ITMX

 * ETMX

 * ITMY

 * ETMY


 * Long term:
  * 7/1-8/9: FPMI/DRFPMI commissioning
  * 8/12-10/4: ASC
  * 10/7-11/15: Post Commissioning
  * 11/18-12/13: Noise hunting


==== Note ====

==== Night work ====

==== Weekend ====

==== Report of work ====
 * DGS maintenance
  * 2 BIO cards installed to k1als0.
   * cable will be connected today.
  * Replaced RT-PCs for PR2 and PR3.
 * OMC cavity was bad, and the transmission was measured as ~20%. Finesse was roughly ~300.
  * OMC REFL was blocked and not coming instead of just an AR reflection.
  * Mirror alignment will be changed.
 * 0.5C temperature shift happened in the center room.
 * SR3 H2 noise was temporarily fixed using spare ADC port.
  * Broken LVDT driver was exchanged and to be fixed for tuning.
 * Tried X arm transition but IMC did not work well.

==== Work plan ====